OpenStudy (squishy):

Solve using substituion.Give solutions in alphabetical order. y = 4x + 2 y = 6x - 10 Need answer and explanation.Thanks

OpenStudy (anonymous):

y=4x+2, substitute it into the second equation: 4x+2=6x-10 12=2x x=6 plug this into either equation to find y y=24+2=26 x=6,y=26

OpenStudy (squishy):

Thanks so much